# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 7.3. KSBA stability of the mirror family
By Lemma 7.11, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$. Hence, $\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{Y}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$, compatibly with the theta functions. In particular, it follows from the above description of $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ in terms of theta functions that $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{D}_{\mathcal{X}_{\sigma}}$, and, by Equation (6.6), $\mathcal{C}_{\sigma}:=\mathcal{C}_{\mathcal{X}}|_{\mathcal{Y}_{\sigma}}$ is the base change of $\mathcal{C}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$.
Finally, the map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) is a dominant toric morphism and so restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. By Theorem 7.7 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the restriction $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$ is irreducible. Hence, the restriction $\mathcal{Y}_{\sigma}|_{T_{H}}$ is also irreducible, and we conclude that $\mathcal{Y}_{\sigma}$ is an irreducible component of $\mathcal{Y}$. ∎
@@
Let $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ be a quasi-projective open Kulikov degeneration. Then, $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is a family of KSBA stable pairs.
@@
For a torus orbit corresponding to a cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$, we can assume, up to replacing $\mathcal{X}$ by another projective crepant resolution of $\overline{\mathcal{X}}$, that this cone is a face $H$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$. By the description of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$ given in Theorem 7.12 i), and by [50, Proposition 7.2], it is enough to show that for every 2-dimensional face $\sigma$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$, the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a family of KSBA stable pairs. As the map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) is a dominant toric morphism, it restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. Hence, by Theorem 7.12 ii), the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a base change of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{X_{\sigma}})$ restricted to $T_{\mathcal{X}_{\sigma}}$. By Theorem 7.7 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the later family is a family of KSBA stable pairs and this concludes the proof in this case.
@@
For a general torus orbit, we can assume as in Remark 6.29 that the corresponding cone is a bogus cone of the form $H+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$, with $H$ as above and $I$ a subset $P_{\mathbb{Z}}$ consisting of integral points $v$ such that $X^{v}$ is contracted by $f_{H}$. By Theorem 6.28, the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to such a torus orbit is obtained by $T_{0}$-equivariance from the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to $T_{H}$. As the base change in Theorem 7.12 is compatible with the $T_{0}$-actions, it is enough to show that, for every 2-dimensional face $\sigma$ of $\mathscr{P}_{H}$, the extension of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the toric stratum corresponding to the bogus cone $\sum_{v\in I_{\sigma}}\mathbb{R}_{\geq 0}[X^{v}_{\sigma}]$ of $\mathcal{X}_{\sigma}$, where $I_{\sigma}=I\cap\sigma_{\mathbb{Z}}$, is a family of KSBA stable pairs. By Theorem 6.28, any fiber $Y_{\sigma,t}$ of this family is related by torus action, and so is isomorphic, to a fiber of $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$. By Theorem 7.7 ii), such a fiber is normal, and so the KSBA stability follows by Theorem 7.7 i). ∎

### v3

@@ 7.2. KSBA stability of the mirror family
By Lemma 7.5, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})$. Hence, $\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{Y}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$, compatibly with the theta functions. In particular, it follows from the above description of $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ in terms of theta functions that $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ is the base change of $\mathcal{D}_{\mathcal{X}_{\sigma}}$, and, by Equation (6.8), $\mathcal{C}_{\sigma}:=\mathcal{C}_{\mathcal{X}}|_{\mathcal{Y}_{\sigma}}$ is the base change of $\mathcal{C}_{\mathcal{X}_{\sigma}}$ along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$. The toric morphism $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$ between the dense torus orbits. By Theorem 7.1 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the restriction $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$ is irreducible. Hence, the restriction $\mathcal{Y}_{\sigma}|_{T_{H}}$ is also irreducible, and we conclude that $\mathcal{Y}_{\sigma}$ is an irreducible component of $\mathcal{Y}$. ∎
@@
Let $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ be a generic quasi-projective open Kulikov degeneration. Then, the mirror family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is a family of KSBA stable pairs.
@@
For a torus orbit corresponding to a cone of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, we can assume, up to replacing $\mathcal{X}$ by another projective crepant resolution of $\mathcal{X}^{\mathrm{can}}$, that this cone is a face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. By the description of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$ given in Theorem 7.6 i), and by [53, Proposition 7.2], it suffices to show that for every 2-dimensional face $\sigma$ of $\mathscr{P}_{H}$, the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a family of KSBA stable pairs. The toric morphism $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ in Equation (7.1) restricts to a morphism $T_{H}\rightarrow T_{\mathcal{X}_{\sigma}}$. Hence, by Theorem 7.6 ii), the restriction of $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})$ to $T_{H}$ is a base change of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{X_{\sigma}})$ restricted to $T_{\mathcal{X}_{\sigma}}$. By Theorem 7.1 ii) applied to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$, the latter family is a family of KSBA stable pairs, and this concludes the proof in this case.
@@
For a general torus orbit, we can assume, as in Remark 6.42, that the corresponding cone is a bogus cone of the form $H+\sum_{v\in I}\mathbb{R}_{\geq 0}[X^{v}]$, where $H$ is as described above and $I$ a subset $P_{\mathbb{Z}}$ consisting of integral points $v$ such that $X^{v}$ is contracted by $f_{H}$. By Theorem 6.41, the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to such a torus orbit is obtained by $T_{0}$-equivariance from the restriction of $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ to $T_{H}$. Since the base change in Theorem 7.6 is compatible with the $T_{0}$-actions, it suffices to show that, for every 2-dimensional face $\sigma$ of $\mathscr{P}_{H}$, the extension of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the toric stratum corresponding to the bogus cone $\sum_{v\in I_{\sigma}}\mathbb{R}_{\geq 0}[X^{v}_{\sigma}]$ of $\mathcal{X}_{\sigma}$, where $I_{\sigma}=I\cap\sigma_{\mathbb{Z}}$, is a family of KSBA stable pairs. By Theorem 6.41, any fiber $Y_{\sigma,t}$ of this family is related by torus action, and so is isomorphic, to a fiber of $\mathcal{Y}_{\mathcal{X}_{\sigma}}|_{T_{\mathcal{X}_{\sigma}}}$. By Theorem 7.1 ii), such a fiber is normal, and so the KSBA stability follows by Theorem 7.1 i). ∎
7.3. Non-constancy of the mirror family
In this section, we prove that the mirror family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is non-constant in restriction to one-dimensional toric strata of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$.
Theorem 7.8.
Let $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ be a generic quasi-projective open Kulikov degeneration. Then, the family of KSBA stable pairs $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\rightarrow\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ is non-constant in restriction to every one-dimensional toric stratum of the toric variety $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$.
Proof.
Let $\rho_{0}$ be a codimension one cone of the secondary fan $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, corresponding to a 1-dimensional toric stratum $S_{\rho_{0}}$ of $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. As the fan $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ is complete, there are two maximal cones $\rho_{1}$ and $\rho_{2}$ containing $\rho_{0}$, corresponding to the two 0-dimensional toric strata $S_{\rho_{1}}$ and $S_{\rho_{2}}$ contained in $S_{\rho_{0}}$. We will show that at least one of the fibers $(Y_{\rho_{1}},D_{\rho_{1}})$ or $(Y_{\rho_{2}},D_{\rho_{2}})$ of $(\mathcal{Y},\mathcal{D})$ over the points $S_{\rho_{1}}$ and $S_{\rho_{2}}$ is not isomorphic to the general fiber $(Y_{\rho_{0}},D_{\rho_{0}})$ of $(\mathcal{Y},\mathcal{D})$ over $S_{\rho_{0}}$.
Let $\tilde{\rho}_{0}$ be a codimension one cone of the Mori fan containing $\rho_{0}$, and $\tilde{\rho}_{1}$, $\tilde{\rho}_{2}$ be the two maximal dimensional cones of the Mori fan containing $\tilde{\rho}_{0}$ and contained in $\rho_{1}$ and $\rho_{2}$ respectively. For $i=0,1,2$, we have $\tilde{\rho}_{i}=\mathrm{Nef}(\mathcal{Z}_{i}/\mathcal{X}^{\mathrm{can}})+\sum_{v\in J_{i}}\mathbb{R}_{\geq 0}[X^{v}]$, where $\mathcal{Z}_{i}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a partial crepant resolution, and $X^{v}$ with $v\in J_{i}$ are divisors which are contracted by a crepant resolution of $\mathcal{Z}_{i}$. Moreover, since $\tilde{\rho}_{0}$ is a codimension one face of $\tilde{\rho}_{i}$, there exists a corresponding birational contraction $\mathcal{Z}_{i}\rightarrow\mathcal{Z}_{0}$. This map is either of relative Picard rank one, in which case $J_{i}=J_{0}$, or it is an isomorphism, in which case $J_{i}=J_{0}\sqcup\{V\}$ for some $V\in J_{i}\setminus J_{0}$. The latter case cannot happen for both $i=1$ and $i=2$: indeed, if $J_{1}=J_{0}\sqcup\{V\}$, then $X^{V}]\in\tilde{\rho}_{1}$ and $X^{V}]\notin\tilde{\rho}_{0}$, and so $X^{V}]\notin\tilde{\rho}_{2}$.
Hence, without loss of generality, we can assume that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is of relative Picard rank one and $J_{1}=J_{0}$. Let $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{1}$ be a projective crepant resolution, and denote by $H_{1}$ the face of $NE(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ corresponding to $\mathcal{X}^{\prime}\rightarrow\mathcal{Z}_{1}$, and by $H_{0}$ the codimension one face of $H_{1}$ corresponding to $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$. By Theorem 7.6, the intersection complex of $(Y_{\rho_{1}},D_{\rho_{1}})$ (resp. $(Y_{\rho_{0}},D_{\rho_{0}})$) is given by the polyhedral decomposition $\mathscr{P}_{H_{1}}$ (resp. $\mathscr{P}_{H_{0}}$) associated to $H_{1}$ (resp . $H_{0}$) as in §7.2. We divide the remainder of the proof into two cases, depending if $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is an isomorphism in codimension one or not.
Case 1: Assume that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is an isomorphism in codimension one. If $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ contracts a curve in the double locus of $\mathcal{Z}_{1}$, then the corresponding edge in $\mathscr{P}_{H_{1}}$ disappears in $\mathscr{P}_{H_{0}}$. Hence, $Y_{\rho_{1}}$, and $Y_{\rho_{0}}$ are not isomorphic, and the result is proved in this case. On the other hand, if $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ does not contract any curve in the double locus of $\mathcal{Z}_{1}$, then the exceptional locus of $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is a rigid interior curve in an irreducible component of $\mathcal{Z}_{1}$, whose strict transform in $\mathcal{X}^{\prime}$ is necessarily an interior exceptional curve. It follows that $NE(\mathcal{Z}_{0}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$ is contained in the codimension one face of $NE(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$ defined by a multiple M1 flop contracting this interior exceptional curve, which is by Lemma 6.38 in contradiction with the assumption that $\rho_{0}$ is a codimension cone of the secondary fan, and so this last case actually does not occur.
Case 2: Assume that $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is a divisorial contraction. Since $\mathcal{Z}_{1}\rightarrow\mathcal{Z}_{0}$ is of relative Picard rank one, its exceptional locus consists of a single divisor $Z_{1}^{v}$ corresponding to a vertex $v$ of $\mathscr{P}_{H_{1}}$. If an irreducible component of the boundary of $Z_{1}^{v}$ is contracted into $\mathcal{Z}_{0}$, then the corresponding edge of $\mathscr{P}_{H_{1}}$ disappears in $\mathscr{P}_{H_{0}}$, and so we also deduce in this case that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are not isomorphic. If no boundary component of $Z_{1}^{v}$ is contracted in $\mathcal{Z}_{0}$, then, as in [53, §8.2], $Z_{1}^{v}$ is necessarily a generically $\mathbb{P}^{1}$-bundle over $\mathbb{P}^{1}$, contracted to its base $\mathbb{P}^{1}$ in $\mathcal{Z}_{0}$, and its boundary $\partial Z_{v}^{1}$ is either an irreducible nodal curve mapping 2:1 to the base, or the union of two sections intersecting in two points.
If $\partial Z_{v}^{1}$ is the union of two irreducible components $D_{1}$ and $D_{2}$, consider the corresponding edges $E_{1}$ and $E_{2}$ adjacent to $v$ in $\mathscr{P}_{H_{1}}$, and choose a 2-dimensional face $\sigma_{1}$ of $\mathscr{P}_{H_{1}}$ such that $E_{1}\cup E_{2}$ are edges of the boundary of $\sigma_{1}$. Denote by $\sigma_{0}$ the same face viewed as a 2-dimensional face of $\mathscr{P}_{H_{0}}$. By Theorem 7.6, there is an irreducible component $(Y_{\sigma_{1}},\partial Y_{\sigma_{1}})$ (resp. $(Y_{\sigma_{0}},\partial Y_{\sigma_{0}})$) of $Y_{\rho_{1}}$ (resp. $Y_{\rho_{0}}$) isomorphic to the general fiber of the mirror family to the open Kulikov degeneration $\mathcal{X}_{\sigma_{1}}$ (resp. $\mathcal{X}_{\sigma_{0}}$) with tropicalization $(C(\sigma_{1}),C(\mathscr{P}_{\sigma_{1}}))$ (resp. $(C(\sigma_{0}),C(\mathscr{P}_{\sigma_{0}}))$). The vertex $v$ can be viewed as a boundary vertex of both $\mathscr{P}_{\sigma_{1}}$ and $\mathscr{P}_{\sigma_{0}}$. Since the $\mathbb{P}^{1}$-fibers of $Z^{1}_{v}$ are not contracted in $\mathcal{Z}_{1}$, the non-compact irreducible component corresponding to $v$ in the central fiber $\mathcal{X}_{\sigma_{1},0}$ of $\mathcal{X}_{\sigma_{1}}$ is a 0-surface. Hence, by Lemma 6.35, it corresponds to a singular point of $\partial Y_{\sigma_{1}}$. On the other hand, since the $\mathbb{P}^{1}$-fibers of $Z^{1}_{v}$ are contracted in $\mathcal{Z}_{0}$, the non-compact irreducible component corresponding to $v$ in the central fiber $\mathcal{X}_{\sigma_{0},0}$ of $\mathcal{X}_{\sigma_{0}}$ is a $1$-surface. Thus, by Lemma 6.35, it corresponds to a smooth point of $\partial Y_{\sigma_{0}}$. Since a singular point of $\partial Y_{\sigma_{1}}$ becomes a smooth point in $\partial Y_{\sigma_{0}}$, it follows that $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are not isomorphic.
Finally, the case where $\partial Z^{1}_{v}$ is irreducible follows from the previous case as in [53, §8.2]. Indeed, denoting by $\sigma_{1}$ (resp. $\sigma_{0}$) the 2-dimensional face of $\mathscr{P}_{H_{1}}$ (resp. $\mathscr{P}_{H_{0}}$) containing $v$, and cutting $\sigma_{1}$ (resp. $\sigma_{0}$) along the edges of $\mathscr{P}_{\sigma_{1}}$ (resp. $\mathscr{P}_{\sigma_{0}}$) connecting $v$ to the boundary of $\sigma_{1}$ (resp. $\sigma_{0}$), we obtain an integral affine manifold with singularities $(\tilde{\sigma}_{1},\mathscr{P}_{\tilde{\sigma}_{1}})$ (resp. $(\tilde{\sigma}_{0},\mathscr{P}_{\tilde{\sigma}_{0}})$) with the same formal properties as $(\sigma_{1},\mathscr{P}_{\sigma_{1}})$ (resp. $(\sigma_{0},\mathscr{P}_{\sigma_{0}})$) in the case where $\partial Z^{1}_{v}$ is reducible. There is a natural open Kulikov degeneration $\mathcal{X}_{\widetilde{\sigma}_{1}}$ (resp. $\mathcal{X}_{\widetilde{\sigma}_{0}}$) with tropicalization $(C(\widetilde{\sigma}_{1}),C(\mathscr{P}_{\widetilde{\sigma}_{1}}))$ (resp. $(C(\widetilde{\sigma}_{0}),C(\mathscr{P}_{\widetilde{\sigma}_{0}}))$), and the irreducible component $Y_{\tilde{\sigma}_{1}}$ (resp. $Y_{\tilde{\sigma}_{0}}$) of the general fiber of the corresponding mirror family is the normalization of $Y_{\sigma_{1}}$ (resp. $Y_{\sigma_{0}}$). As above, a singular point of $\partial Y_{\tilde{\sigma}_{1}}$ becomes smooth in $\partial Y_{\tilde{\sigma}_{0}}$. Consequently, $Y_{\rho_{1}}$ and $Y_{\rho_{0}}$ are also non-isomorphic in this case. ∎